OpenStudy (anonymous):

Is the number of games won by a major league baseball team in a season related to the team batting average? The table below shows the number of games won and the batting average of 8 teams. Team - Games Won - Batting Average 1- 115- 0.285 2- 66- 0.267 3- 66- 0.263 4- 103- 0.274 5- 62- 0.26 6- 76- 0.282 7- 71- 0.266 8- 110- 0.282 Using games won as the independent variable x, do the following: (a) The correlation coefficient is r = _________________ (b) The equation of the least squares line is y = _________ + _____________ x 1 hour ago - 4 days left to answer.

OpenStudy (tkhunny):

What's your plan? Won't you need \(\sum x\;and\;\sum y\;and\; \sum xy\)? That's a good place to start.

OpenStudy (anonymous):

That's the thing. I'm not sure how to get those. I can't seem to get any closer to solve this. Any help with how to approach this is appreciated.

OpenStudy (tkhunny):

What do you mean you can't get those. Just add them up! 115 + 66 + 66 + 103 + 62 + 76 + 71 + 110 = Sum of the x's

OpenStudy (anonymous):

Okay, I found that x = 669 and y = 2.179 What should my next step be? I'm not sure how to find the correlation coefficient.

OpenStudy (tkhunny):

Sum of the xy. Multiply the two and than add up all those. You will also need the sum of the x^2 and y^2 It can be a little tedious.

OpenStudy (anonymous):

669 * 2.179 = 1457.751 669 + 2.179 = 671.179 669^2 * 2.179^2 = 2125037.978001 I think my numbers should be right.

OpenStudy (tkhunny):

You should get: \(n = 8\) \(\sum x = 2.185\) \(\sum y = 669\) \(\sum xy = 183.807\) \(\sum x^{2} = 0.597319\) \(\sum y^{2} = 59307\) I can't really tell what you are doing. Maybe you are forgetting that \(\sum xy \ne \sum x \cdot \sum y\)

OpenStudy (tkhunny):

Whoops! I'm backwards. Trade x and y. That doesn't actually change anything, except which is independent and which is dependent.

OpenStudy (anonymous):

Oh, I see where I messed up. Thank you for that. I assume my next step should be to plug the numbers in an equation correct?

OpenStudy (tkhunny):

I sincerely hope you already have the equation. I know I don't want to type it all. :-)

OpenStudy (anonymous):

It's this one, correct? Also thank you so much for the help, you're really helping me out of a jam. :)

OpenStudy (tkhunny):

Yes, that's it. How about that least squares equation?

OpenStudy (tkhunny):

You already have all the numbers. Just different equations.

OpenStudy (anonymous):

No, that equation I don't have.

OpenStudy (tkhunny):

Okay, I'll type in the formulas for a and b, given y = ax+b Denominator for Both = \(\left(\sum x \right)^{2} - n\cdot \sum x^{2}\) a-numerator = \(\sum x\sum y - n\sum xy\) b-numerator = \(\sum x \sum xy - \sum x^{2} \sum y\) That's all. Just calculate those two ratios.

OpenStudy (anonymous):

Oh my gosh, thank you so much! You were such a big help!